# A proof of the Kim-Vu sandwich conjecture

### v1

@@ 6.1 Pseudorandom properties for lower bounds for path counts
$|E_{\geq 2,K}(X_{2j-1}\cup Y_{2j-1})|$ $\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{j=\ell_{1}\}}\right)\cdot d|X_{2j-1}\cup Y_{2j-1}|$
$\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{j=\ell_{1}\}}\right)\cdot 3\delta^{j}d^{j}r.$ (41)
Furthermore, we have that $\cup_{a=0}^{2j-1}(X_{a}\cup Y_{a})\subset\left(\cup_{a=0}^{2j-2}(X_{a}\cup Y_{a})\right)\cup N_{F\setminus K}\left(\cup_{a=0}^{2j-2}(X_{a}\cup Y_{a})\right)$, and $|\cup_{a=0}^{2j-2}(X_{a}\cup Y_{a})|\leq 4d^{j-1}\delta^{j-1}r\leq 4|\cup_{a=0}^{2j-1}(X_{a}\cup Y_{a})|/\delta$, while
$\left|\cup_{a=0}^{2j-1}(X_{a}\cup Y_{a})\right|\leq 3d^{j-1}\delta^{j}r\leq\frac{3\lambda n}{2\cdot 10^{7}}\cdot d^{j-1-\ell_{1}}\delta^{j-\ell_{1}}\leq\left\{\begin{array}[]{ll}\frac{\lambda n}{10^{6}d\log n}&\text{ if }j<\ell_{1},\\ \frac{\lambda n}{10^{6}d}&\text{ if }j=\ell_{1}.\end{array}\right$.
@@
$e_{K}\left(\cup_{a=0}^{2j-1}(X_{a}\cup Y_{a})\right)$ $\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{j=\ell_{1}\}}\right)\cdot d\left|\cup_{a=0}^{2j-1}(X_{a}\cup Y_{a})\right|$
$\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{j=\ell_{1}\}}\right)\cdot 3\delta^{j}d^{j}r.$ (42)
@@
$\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{j\in\{1,\ell_{1}\}\}}\right)\cdot 8\cdot d\cdot\sum_{v\in X_{2j-2}}d_{F\setminus K}(v).$ (43)
@@
$\hskip 227.62204pt-(1+\lambda)\delta(d|X_{2j-1}|-|X_{2j}|)$
@@
$|X_{2\ell_{1}}|\geq\frac{\sum_{v\in X_{2\ell_{1}}}d_{F\setminus K}(v)}{(1+\lambda)\delta}\geq\frac{(1-40\lambda)\cdot(1-\lambda)d^{\ell_{1}}\delta^{\ell_{1}+1}}{(1+\lambda)\delta}\geq(1-45\lambda)d^{\ell_{1}}\delta^{\ell_{1}}r\geq\frac{\lambda n}{10^{8}},$
@@
Let $s=|U_{2}|$ and $t=|V_{2}|$, so that $s,t\geq r$, and enumerate $U_{2}=\{u_{1},\ldots,u_{s}\}$ and $V_{2}=\{v_{1},\ldots,v_{t}\}$. For each $i\in[s]$, pick $u^{\prime}_{i}\in N_{K}(u_{i})\cap U_{1}$ and for each $j\in[t]$ pick $v_{j}^{\prime}\in N_{K}(v_{j})\cap V_{1}$. For each $i\in[s]$, let $\tilde{U}_{i}=\{u^{\prime}_{i},u^{\prime}_{i+1},\ldots,u^{\prime}_{i+r-1}\}$ and $\hat{U}_{i}=\{u_{i},u_{i+1},\ldots,u_{i+r-1}\}$, and, for each $j\in[t]$, let $\tilde{V}_{j}=\{v^{\prime}_{j},v^{\prime}_{j+1},\ldots,v^{\prime}_{j+r-1}\}$ and $\hat{V}_{j}=\{v_{j},v_{j+1},\ldots,v_{j+r-1}\}$, with addition in the indices modulo $s$ and $t$ as appropriate. For each $i\in[s]$ and $j\in[t]$, let $W_{i,j}=W\cup\{u,v\}\cup\tilde{U}_{i}\cup\tilde{V}_{j}$, so that $|W_{i,j}|\leq 3r+\log{n}+\lambda\delta$. Now, by Claim 6.4, there are at least $\left(1-\frac{C_{0}\lambda}{4}\right)\cdot\frac{r^{2}\delta^{2\ell_{1}+1}d^{2\ell_{1}}}{n}$ $(F\setminus K,K)$-alternating $\hat{U}_{i},\hat{V}_{j}$-paths with length $4\ell_{1}+1$ with no internal vertices in $W_{i,j}$, each of which can be extended to an $(F\setminus K,K)$-alternating $u,v$-path with length $4\ell_{1}+5$ with no internal vertices in $W$. Therefore, as each such path appears for $r^{2}$ pairs $(i,j)$, the number of $(F\setminus K,K)$-alternating $u,v$-paths with length $4\ell_{1}+5$ with no internal vertices in $W$ is at least
@@
Finally, we can show that LABEL:prop:lowerboundKFKalternating:pseud:newnew holds. Let $x,y\in[n]$ be distinct and let $Z\subset[n]$ with $|Z|\leq 40\log n$. Let $y^{\prime}\in N_{K}(y)\setminus(Z\cup\{x\})$, with $d-|(Z\cup\{x\})\cap N_{K}(y)|$ choices. For each $i\geq 0$, let $\mathcal{P}_{i}$ be the set of pairs $(v,P)$ such that $P$ is an $(F\setminus K,K)$-alternating $y^{\prime},v$-path of length $2i$ with no vertices in $Z\cup\{x\}$ and $d_{F\setminus K}(v,Z\cup V(P)\cup\{x,y\})\leq\left(\mathbf{1}_{\{i=0\}}+1/\log{n}\right)\lambda\delta$. Note that $|\mathcal{P}_{0}|=1$ as $|Z\cup\{x,y\}|\leq\lambda\delta$.
@@
Let $(v,P)\in\mathcal{P}_{i-1}$. Let $Z^{+}=V(P)\cup Z\cup\{x,y\}$, so that $|Z^{+}|\leq 100\log n$. Let $B$ be the set of $w\in[n]\setminus Z^{+}$ with $d_{F\setminus K}(w,Z^{+})\geq 10$ so that, by LABEL:prop:new, we have $|B|\leq 100\log n$.
@@
$\leq\left(4\cdot\mathbf{1}_{\{i=1\}}+10/\log{n}\right)\cdot\lambda\delta^{2}d.$
@@
Then, let $P^{\prime}$ be an $(F\setminus K)$-alternating $x,v$-path of length $4\ell_{1}+5$ with no internal vertices in $Z\cup V(P)\cup\{y\}$, with at least $(1-C_{0}\lambda/2)\delta^{\ell_{1}+3}d^{\ell_{1}+2}/n$ options by Claim 6.5. Then, $(P\cup P^{\prime})+y^{\prime}y$ is an $(F\setminus K,K)$-alternating $x,y$-path with length $2\ell$. Combining our choices for $y^{\prime},P$ and $P^{\prime}$, the number of choices is at least
$(d-|(Z\cup\{x\})\cap N_{K}(y)|)\left(1-\frac{C_{0}\lambda}{4}\right)\delta^{\ell_{2}}d^{\ell_{2}}$ $\cdot\left(1-\frac{C_{0}\lambda}{2}\right)\frac{\delta^{\ell_{1}+3}d^{\ell_{1}+2}}{n}$
@@ 6.2 Likely properties of $G_{d}(n)$: expansion
$\sum_{i=\lambda d\delta m/16}^{n^{2}}\frac{|\mathcal{K}^{i}|}{|\mathcal{K}_{d}(n)|}\leq n^{2}\exp\left(-\frac{Cm\log n}{16}\right).$
@@ 6.3 Likely properties of $F\setminus K$: expansion
Let $1/n\ll 1/C\ll\varepsilon\leq 1$. Let $d$ and $\lambda$ satisfy $\log^{-3}n\leq\lambda\ll 1$, $10^{3}d\leq\lambda n$, and $d\geq C\log n$. Let $m\in\mathbb{N}$ satisfy $\varepsilon dn\leq m\leq n^{1+\sigma}$. Let $\delta=2m/n$ and suppose that $\lambda\delta\geq C\log n$. Let $p=m(\binom{n}{2}-dn/2)^{-1}$. Let $K\sim K_{d}(n)$, form $E$ by including each element of $n]^{(2)}\setminus E(K)$ uniformly at random with probability $p$, and let $F=K+E$.
@@
Let $U\subset[n]$ with $|U|\leq\lambda n/100\delta\log n$, and set $m=|U|$. We will show that, with probability $1-\exp\left(-\omega(m)\right)$, $U$ satisfies the property we want, and then take a union bound over all sets $U$ we must consider.
Let $X=e(F[U])$, so that again, we have
$\mathbb{P}\left(X\geq\frac{\lambda\delta m}{2\log n}\right)$ $\leq\binom{m^{2}}{\frac{\lambda\delta m}{2\log n}}\cdot p^{\frac{\lambda\delta m}{2\log n}}\leq\left(\frac{2em^{2}p\log n}{\lambda\delta m}\right)^{\frac{\lambda\delta m}{2\log n}}\leq\left(\frac{4em\log n}{\lambda n}\right)^{\frac{\lambda\delta m}{2\log n}}.$
@@
$\left(\frac{4em\log n}{\lambda n}\right)^{\frac{\lambda\delta m}{2\log n}}\leq\left(\frac{4e}{100\delta}\right)^{\frac{\lambda\delta m}{2\log n}}\leq\left(\frac{4e}{100\delta}\right)^{\frac{Cm}{2}}=\exp(-\omega(m)),$
as $\delta\geq C\log n$. Moreover, the number of sets $U\subset[n]$ with $|U|=m$ for which there is some $U^{\prime}\subset[n]$ with $|U^{\prime}|\leq 4m/d$ and $U\subset U^{\prime}\cup N_{K}(U^{\prime})$ is (as such a set $U^{\prime}$ has $|N_{K}(U^{\prime})|\leq d|U^{\prime}|\leq 4m$) at most $\binom{n}{4m/d}\cdot 2^{5m}\leq n^{4m/d}\cdot 2^{5m}\leq 2^{6m}$ as $d\geq C\log n$. Note that these conditions imply that if there is some such $U$ then we must have $1\leq 4m/d$, sot that $m\geq d/4\geq C\log n/4$ Thus, the probability that LABEL:eq:fewdoubled:F-K:bothcases does not hold is at most
$\sum_{m=C\log n/4}^{n}2^{6m}\cdot\exp\left(-\omega(m))\right)=o(n^{-14}),$
@@
Let $1/n\ll 1/C\ll\varepsilon,\sigma\ll 1$. Let $d$ and $\lambda$ satisfy $\log^{-3}n\leq\lambda\ll 1$, $10^{3}d\leq\lambda n$, and $d\geq C\log n$. Let $m\in\mathbb{N}$ satisfy $\varepsilon dn\leq m\leq n^{1+\sigma}$. Let $\delta=2m/n$ and suppose that $\lambda\delta\geq C\log n$. Let $p=m(\binom{n}{2}-dn/2)^{-1}$. Let $K\sim K_{d}(n)$, form $E$ by including each element of $n]^{(2)}\setminus E(K)$ uniformly at random with probability $p$, and let $F=K+E$. Then, with probability $1-o(n^{-14})$, the following holds.
@@
Let $U\subset[n]$ with $|U|\leq\lambda n/100\delta\log n$, and set $m=|U|$. We will show that, with probability $1-\exp\left(-\omega(m)\right)$, $U$ satisfies the property we want, so that the result follows by a union bound over all such $U$ for which there is some $U^{\prime}\subset[n]$ with $|U^{\prime}|\leq 4|U|/d$ and $U\subset U^{\prime}\cup N_{K}(U^{\prime})$, as at the end of Lemma 6.12.
Consider a process in which we reveal edges in $G$ by taking each $v\in[n]\setminus U$ and revealing potential edges in $G$ into $U$ until a neighbour of $v$ in $U$ has been revealed or it has been revealed that $v$ has no neighbours in $U$. Let $V$ be the set of vertices $v$ after this which have a neighbour in $U$. Then, reveal the rest of the edges between $V$ and $U$ that have not yet been considered, and let $X$ be the number of edges between $V$ and $U$ revealed here. Note that $|E_{\geq 2,F^{\prime}}(U)|\leq 2X$.
Now, let $\mathcal{E}$ be the event that $|V|\leq 3\delta m$. As $V=N_{F}(U)$, by a simple application of Chernoff’s bound, with probability $1-\exp(-\omega(m))$, $\mathcal{E}$ holds. Let $Y=X$ if $\mathcal{E}$ holds, and $Y=0$ otherwise. Thus, with probability $1-\exp(-\omega(m))$, $|E_{\geq 2,F^{\prime}}(U)|\leq 2Y$.
For $M=\frac{\lambda\delta m}{4\log n}=\Omega(m)$, we have, then
$\mathbb{P}(Y\geq M)$ $\leq\binom{m\cdot 3\delta m}{M}q^{M}\leq\left(\frac{3e\delta m^{2}q}{M}\right)^{M}=\left(\frac{6e\delta^{2}m^{2}}{Mn}\right)^{M}=\left(\frac{24e\delta m}{\lambda n}\right)^{M}\leq\left(\frac{24e}{100\log n}\right)^{M}$
$=\exp(-\omega(m)).$
As $|E_{\geq 2,F^{\prime}}(U)|\leq 2Y$ with probability $1-\exp(-\omega(m))$, we have that LABEL:eq:fewdoubled:F-K:bothcases2 holds for $U$ with probability $1-\exp(-\omega(m))$. Therefore, similarly to the end of the proof of Lemma 6.12, we have that LABEL:eq:fewdoubled:F-K:bothcases2 holds with probability $1-o(n^{-14})$, as required. ∎
@@
Let $1/n\ll 1/C\ll\varepsilon,\sigma\ll 1$. Let $0<p\leq n^{\sigma-1}$ and let $H\sim G(n,p)$. Then, with probability $1-o(n^{-14})$, for each $U\subset[n]$ with $|U|\leq 100\log n$, $|\{v\in[n]\setminus U:d_{H}(v)\geq 10\}|\leq 100\log n$.
@@
Let $U\subset[n]$ satisfy $|U|\leq 100\log n$. For each $v\in[n]\setminus U$, the probability that $d_{H}(v)\geq 10$ is at most $|U|^{10}p^{10}\leq n^{-9}$. Then, the probability that $|\{v\in[n]\setminus U:d_{H}(v)\geq 10\}|>100\log n$ is, setting $k=100\log n$, at most $n^{k}\cdot n^{-9k}k=n^{-8k}$. Taking a union bound over all sets $U\subset[n]$ satisfying $|U|\leq 100\log n$, the property in the lemma holds with probability at most $2n^{k}\cdot n^{-8k}=o(n^{-14})$, as required. ∎

### v2

@@ 6.1 Pseudorandom properties for lower bounds for path counts
$|E_{\geq 2,K}(X_{2i-1}\cup Y_{2i-1})|$ $\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{i=\ell_{1}\}}\right)\cdot d|X_{2i-1}\cup Y_{2i-1}|$
$\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{i=\ell_{1}\}}\right)\cdot 3\delta^{i}d^{i}r.$ (42)
Furthermore, we have that $\cup_{a=0}^{2i-1}(X_{a}\cup Y_{a})\subset\left(\cup_{a=0}^{2i-2}(X_{a}\cup Y_{a})\right)\cup N_{F\setminus K}\left(\cup_{a=0}^{2i-2}(X_{a}\cup Y_{a})\right)$, and $|\cup_{a=0}^{2i-2}(X_{a}\cup Y_{a})|\leq 3d^{i-1}\delta^{i-1}r\leq 4|\cup_{a=0}^{2i-1}(X_{a}\cup Y_{a})|/\delta$, while
$\left|\cup_{a=0}^{2i-1}(X_{a}\cup Y_{a})\right|\leq 3d^{i-1}\delta^{i}r\leq\frac{3\lambda n}{2\cdot 10^{7}}\cdot d^{i-1-\ell_{1}}\delta^{i-\ell_{1}}\leq\left\{\begin{array}[]{ll}\frac{\lambda n}{10^{6}d\log n}&\text{ if }i<\ell_{1},\\ \frac{\lambda n}{10^{6}d}&\text{ if }i=\ell_{1}.\end{array}\right$.
@@
$e_{K}\left(\cup_{a=0}^{2i-1}(X_{a}\cup Y_{a})\right)$ $\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{i=\ell_{1}\}}\right)\cdot d\left|\cup_{a=0}^{2i-1}(X_{a}\cup Y_{a})\right|$
$\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{i=\ell_{1}\}}\right)\cdot 3\delta^{i}d^{i}r.$ (43)
@@
$\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{i\in\{1,\ell_{1}\}\}}\right)\cdot 8\cdot d\cdot\sum_{v\in X_{2i-2}}d_{F\setminus K}(v).$ (44)
@@
$\hskip 227.62204pt-(1+\lambda)\delta(d|X_{2i-1}|-|X_{2i}|)$
@@
$|X_{2\ell_{1}}|\geq\frac{\sum_{v\in X_{2\ell_{1}}}d_{F\setminus K}(v)}{(1+\lambda)\delta}\geq\frac{(1-40\lambda)\cdot(1-\lambda)d^{\ell_{1}}\delta^{\ell_{1}+1}r}{(1+\lambda)\delta}\geq(1-45\lambda)d^{\ell_{1}}\delta^{\ell_{1}}r\geq\frac{\lambda n}{10^{8}},$
@@
Let $s=|U_{2}|$ and $t=|V_{2}|$, so that $s,t\geq r$, and enumerate $U_{2}=\{u_{1},\ldots,u_{s}\}$ and $V_{2}=\{v_{1},\ldots,v_{t}\}$. For each $i\in[s]$, pick $u^{\prime}_{i}\in N_{K}(u_{i})\cap U_{1}$ and for each $j\in[t]$ pick $v_{j}^{\prime}\in N_{K}(v_{j})\cap V_{1}$. For each $i\in[s]$, let $\tilde{U}_{i}=\{u^{\prime}_{i},u^{\prime}_{i+1},\ldots,u^{\prime}_{i+r-1}\}$ and $\hat{U}_{i}=\{u_{i},u_{i+1},\ldots,u_{i+r-1}\}$, and, for each $j\in[t]$, let $\tilde{V}_{j}=\{v^{\prime}_{j},v^{\prime}_{j+1},\ldots,v^{\prime}_{j+r-1}\}$ and $\hat{V}_{j}=\{v_{j},v_{j+1},\ldots,v_{j+r-1}\}$, with addition in the indices modulo $s$ and $t$ as appropriate. For each $i\in[s]$ and $j\in[t]$, let $W_{i,j}=W\cup\{u,v\}\cup\tilde{U}_{i}\cup\tilde{V}_{j}$, so that $|W_{i,j}|\leq 3r+\lambda\delta$. Now, by Claim 6.4, there are at least $\left(1-\frac{C_{0}\lambda}{4}\right)\cdot\frac{r^{2}\delta^{2\ell_{1}+1}d^{2\ell_{1}}}{n}$ $(F\setminus K,K)$-alternating $\hat{U}_{i},\hat{V}_{j}$-paths with length $4\ell_{1}+1$ with no internal vertices in $W_{i,j}$, each of which can be extended to an $(F\setminus K,K)$-alternating $u,v$-path with length $4\ell_{1}+5$ with no internal vertices in $W$. Therefore, as each such path appears for $r^{2}$ pairs $(i,j)$, the number of $(F\setminus K,K)$-alternating $u,v$-paths with length $4\ell_{1}+5$ with no internal vertices in $W$ is at least
@@
Finally, we can show that LABEL:prop:lowerboundKFKalternating:pseud:newnew holds. Let $x,y\in[n]$ be distinct and let $Z\subset[n]$ with $|Z|\leq 40\log n$. Let $y^{\prime}\in N_{K}(y)\setminus(Z\cup\{x\})$, with $d-|(Z\cup\{x\})\cap N_{K}(y)|$ choices. For each $i\geq 0$, let $\mathcal{P}_{i}$ be the set of pairs $(v,P)$ such that $P$ is an $(F\setminus K,K)$-alternating $y^{\prime},v$-path of length $2i$ with no vertices in $Z\cup\{x\}$ and $|N_{F\setminus K}(v)\cap(Z\cup V(P)\cup\{x,y\})|\leq\left(\mathbf{1}_{\{i=0\}}+1/\log{n}\right)\lambda\delta$. Note that $|\mathcal{P}_{0}|=1$ as $|Z\cup\{x,y\}|\leq\lambda\delta$.
@@
Let $(v,P)\in\mathcal{P}_{i-1}$. Let $Z^{+}=V(P)\cup Z\cup\{x,y\}$, so that $|Z^{+}|\leq 100\log n$. Let $B$ be the set of $w\in[n]\setminus Z^{+}$ with $|N_{F\setminus K}(w)\cap Z^{+}|\geq 10$ so that, by LABEL:prop:new, we have $|B|\leq 100\log n$.
@@
$\leq\left(2\cdot\mathbf{1}_{\{i=1\}}+10/\log{n}\right)\cdot\lambda\delta^{2}d.$
@@
Then, let $P^{\prime}$ be an $(F\setminus K,K)$-alternating $x,v$-path of length $4\ell_{1}+5$ with no internal vertices in $Z\cup V(P)\cup\{y\}$, with at least $(1-C_{0}\lambda/2)\delta^{2\ell_{1}+3}d^{2\ell_{1}+2}/n$ options by Claim 6.5. Then, $(P\cup P^{\prime})+y^{\prime}y$ is an $(F\setminus K,K)$-alternating $x,y$-path with length $2\ell$. Combining our choices for $y^{\prime},P$ and $P^{\prime}$, the number of choices is at least
$(d-|(Z\cup\{x\})\cap N_{K}(y)|)\left(1-\frac{C_{0}\lambda}{4}\right)\delta^{\ell_{2}}d^{\ell_{2}}$ $\cdot\left(1-\frac{C_{0}\lambda}{2}\right)\frac{\delta^{2\ell_{1}+3}d^{2\ell_{1}+2}}{n}$
@@ 6.2 Likely properties of $G_{d}(n)$: expansion
$\sum_{i=\lambda d\delta m/16}^{n^{2}}\frac{|\mathcal{K}^{i}|}{|\mathcal{K}_{d}(n)|}\leq n^{2}\exp\left(-\frac{Cm\log n}{64}\right).$
@@ 6.3 Likely properties of $F\setminus K$: expansion
Let $1/n\ll 1/C\ll\varepsilon\leq 1$. Let $d$ and $\lambda$ satisfy $\log^{-3}n\leq\lambda\ll 1$, $10^{3}d\leq\lambda n$, and $d\geq C\log n$. Let $m\in\mathbb{N}$ satisfy $\varepsilon dn/2\leq m\leq n^{1+\sigma}$. Let $\delta=2m/n$ and suppose that $\lambda\delta\geq C\log n$. Let $p=m(\binom{n}{2}-dn/2)^{-1}$. Let $K\sim G_{d}(n)$, form $E$ by including each element of $n]^{(2)}\setminus E(K)$ uniformly at random with probability $p$, and let $F=K+E$.
@@
Let $U\subset[n]$ with $|U|\leq\lambda n/100\delta\log n$ and let $s=|U|$. We will show that, with probability $1-\exp\left(-\omega(s)\right)$, $U$ satisfies the property we want, and then take a union bound over all sets $U$ we must consider.
Let $X=e(F\setminus K[U])$ and let $M=\left\lceil\frac{\lambda\delta s}{2\log{n}}\right\rceil$ so that again, we have
$\mathbb{P}\left(X\geq\frac{\lambda\delta s}{2\log n}\right)\leq\sum_{i=M}^{\binom{s}{2}}\binom{\binom{s}{2}}{i}\cdot p^{i}\leq\sum_{i=M}^{s^{2}}\left(\frac{es^{2}p}{2i}\right)^{i}\leq\sum_{i=M}^{s^{2}}\left(\frac{es^{2}p}{2M}\right)^{i}$
@@
$\frac{es^{2}p}{2M}\leq\frac{esp\cdot\log{n}}{\lambda\delta}\leq\frac{2es\cdot\log{n}}{\lambda n}\leq\frac{2e}{100\delta}$
and $\lambda\delta\geq C\log{n}$ , we thus obtain that
$\mathbb{P}\left(X\geq\frac{\lambda\delta s}{2\log n}\right)\leq\sum_{i=M}^{s^{2}}\left(\frac{2e}{100\delta}\right)^{i}\leq 2\left(\frac{2e}{100\delta}\right)^{\frac{\lambda\delta s}{2\log{n}}}\leq 2\left(\frac{2e}{100\delta}\right)^{\frac{Cs}{2}}=\exp{(-\omega(s))},$
where we have used that $\delta\geq C\varepsilon\log{n}$. Let us now count the number of sets $U\subset[n]$ with $|U|=s$ for which there is some $U^{\prime}\subset[n]$ with $|U^{\prime}|\leq 4s/d$ and $U\subset U^{\prime}\cup N_{K}(U^{\prime})$. Note that the condition $|U^{\prime}|\leq 4|U|/d$ implies that if there is some such $U$ then we must have $1\leq 4s/d$, which in turn implies that $s\geq d/4\geq C\log n/4$. Since $|U^{\prime}\cup N_{K}(U^{\prime})|\leq(d+1)|U^{\prime}|$, the number of possible choices for such a $U\subset U^{\prime}\cup N_{K}(U^{\prime})$ is at most
$\sum_{i=1}^{4s/d}\binom{n}{i}\cdot 2^{(d+1)i}\leq\sum_{i=1}^{4s/d}\left(n2^{(d+1)}\right)^{i}\leq 2\left(n2^{(d+1)}\right)^{4s/d}\leq 2^{5s}$
as $d\geq C\log n$. Thus, the probability that LABEL:eq:fewdoubled:F-K:bothcases does not hold is at most
$\sum_{s=C\log n/4}^{n}2^{5s}\cdot\exp\left(-\omega(s)\right)=o(n^{-14}),$
@@
Let $1/n\ll 1/C\ll\varepsilon,\sigma\ll 1$. Let $d$ and $\lambda$ satisfy $\log^{-3}n\leq\lambda\ll 1$, $10^{3}d\leq\lambda n$, and $d\geq C\log n$. Let $m\in\mathbb{N}$ satisfy $\varepsilon dn/2\leq m\leq n^{1+\sigma}$. Let $\delta=2m/n$ and suppose that $\lambda\delta\geq C\log n$. Let $p=m(\binom{n}{2}-dn/2)^{-1}$. Let $K\sim G_{d}(n)$, form $E$ by including each element of $n]^{(2)}\setminus E(K)$ uniformly at random with probability $p$, and let $F=K+E$. Then, with probability $1-o(n^{-14})$, the following holds.
@@
Let $U\subset[n]$ with $|U|\leq\lambda n/100\delta\log n$, and set $s=|U|$. We will show that, with probability $1-\exp\left(-\Omega(s)\right)$, $U$ satisfies the property we want, so that the result follows by a union bound over all such $U$ for which there is some $U^{\prime}\subset[n]$ with $|U^{\prime}|\leq 4|U|/d$ and $U\subset U^{\prime}\cup N_{K}(U^{\prime})$, as at the end of Lemma 6.12.
Consider a process in which we reveal edges in $E$ by taking each $v\in[n]\setminus U$ and revealing potential edges in $E$ into $U$ until a neighbour of $v$ in $U$ has been revealed or it has been revealed that $v$ has no neighbours in $U$. Let $V$ be the set of vertices $v$ after this which have a neighbour in $U$. Then, reveal the rest of the edges between $V$ and $U$ that have not yet been considered, and let $X$ be the number of edges between $V$ and $U$ revealed here. Note that $|E_{\geq 2,F\setminus K}(U)|\leq 2X$.
Now, let $\mathcal{E}$ be the event that $|V|\leq 3\delta s$. As $V=N_{F\setminus K}(U)$, by a simple application of Chernoff’s bound, with probability $1-\exp(-\omega(s))$, $\mathcal{E}$ holds. Let $Y=X$ if $\mathcal{E}$ holds, and $Y=0$ otherwise. Thus, with probability $1-\exp(-\omega(s))$, $|E_{\geq 2,F\setminus K}(U)|\leq 2Y$.
For $M=\left\lceil\frac{\lambda\delta s}{4\log n}\right\rceil=\Omega(s)$, we have, then
$\mathbb{P}\left(Y\geq\frac{\lambda\delta s}{4\log n}\right)$ $\leq\sum_{i=M}^{|U||V|}\binom{|U||V|}{i}p^{i}\leq\sum_{i=M}^{3\delta s^{2}}\binom{s\cdot 3\delta s}{i}p^{i}\leq\sum_{i=M}^{3\delta s^{2}}\left(\frac{3e\delta s^{2}p}{i}\right)^{i}\leq\sum_{i=M}^{3\delta s^{2}}\left(\frac{3e\delta s^{2}p}{M}\right)^{i}$
As
$\frac{3e\delta s^{2}p}{M}\leq\frac{3esp\cdot 4\log{n}}{\lambda}\leq\frac{24es\delta\log{n}}{\lambda n}\leq\frac{24e}{100}\leq e^{-1/3}$
we see that
$\mathbb{P}\left(Y\geq\frac{\lambda\delta s}{4\log n}\right)\leq\sum_{i=M}^{3\delta s^{2}}e^{-i/3}\leq 4e^{-\frac{\lambda\delta s}{12\log n}}\leq\exp{(-Cs/20)}.$
Thus $|E_{\geq 2,F^{\prime}}(U)|\leq 2Y$ with probability $1-\exp{(-Cs/20)}$. Similarly to the end of the proof of Lemma 6.12, we know that there are at most $2^{5s}$ sets $U$ such that $|U|=s$, there is some set $U^{\prime}$ with $|U^{\prime}|\leq 4s/d$ and $U\subset U^{\prime}\cup N_{K}(U^{\prime})$; and this is only possible is $s\geq d/4\geq C\log{n}/4$. Thus the probability that LABEL:eq:fewdoubled:F-K:bothcases2 does not hold is at most
$\sum_{s=C\log{n}/4}^{n}2^{5s}\cdot\exp{(-Cs/20)}=o(n^{-14})$
as required. ∎
@@
Let $1/n\ll 1/C\ll\varepsilon,\sigma\ll 1$. Let $0<p\leq n^{\sigma-1}$ and let $H\sim G(n,p)$. Then, with probability $1-o(n^{-14})$, for each $U\subset[n]$ with $|U|\leq 100\log n$, $|\{v\in[n]\setminus U:|N_{H}(v)\cap U|\geq 10\}|\leq 100\log n$.
@@
Let $U\subset[n]$ satisfy $|U|\leq 100\log n$. For each $v\in[n]\setminus U$, the probability that $|N_{H}(v)\cap U|\geq 10$ is at most $\binom{|U|}{10}p^{10}\leq|U|^{10}p^{10}\leq n^{-9}$. Then, the probability that $|\{v\in[n]\setminus U:d_{H}(v)\geq 10\}|>100\log n$ is, setting $k=100\log n$, at most $\sum_{i=k}^{n}\binom{n}{i}n^{-9i}\leq\sum_{i=k}^{n}n^{i}n^{-9i}\leq n\cdot n^{-8k}$. Taking a union bound over all sets $U\subset[n]$ satisfying $|U|\leq 100\log n$, the property in the lemma does not hold with probability at most $2n^{k}\cdot n^{1-8k}=o(n^{-14})$, as required. ∎